OpenStudy (anonymous):

Solve r*(1-c*d)=r*(1-c*(d-p))-f for f

OpenStudy (lalaly):

f=r*(1-c*d)-r*(1-c*(d-p))

OpenStudy (anonymous):

try this r-rcd=r*(1-cd-cp)-f r-rcd=r-rcd-rcp-f r-rcd will cancel both sides 0=-rcp-f f=-rcp